A machine of velocity ratio 5 requires 1000j of work to raise aload of 450N through a vertical distance 2m. calculate the efficiency and mechanical advantage of the machine
1
Expert's answer
2013-01-16T12:05:45-0500
Work required Wr = F * d = 450 * 2 = 900 J. Work done Wd =1000 J.
Efficiency E = Wr / Wd = 900 / 1000 = 90%.
Mechanical advantage MA = Vr * E = 5 * 0.9 = 4.5.